# A ‘Catch, Tag, and Release’ Mechanism for Embeddings

### v1

@@ Abstract
Two prominent features of large language models (LLMs) is the presence of large-norm (outlier) features and the tendency for tokens to attend very strongly to a select few tokens. Despite often having no semantic relevance, these select tokens, called attention sinks, along with the large outlier features, have proven important for model performance, compression, and streaming. Consequently, investigating the roles of these phenomena within models and exploring how they might manifest in the model parameters has become an area of active interest. Through an empirical investigation, we demonstrate that attention sinks utilize outlier features to: catch a sequence of tokens, tag the captured tokens by applying a common perturbation, and then release the tokens back into the residual stream, where the tagged tokens are eventually retrieved. We prove that simple tasks, like averaging, necessitate the ‘catch, tag, release’ mechanism hence explaining why it would arise organically in modern LLMs. Our experiments also show that the creation of attention sinks can be completely captured in the model parameters using low-rank matrices, which has important implications for model compression and substantiates the success of recent approaches that incorporate a low-rank term to offset performance degradation. Link to project page.
Attention Sink, Outlier Features, LLMs, Pruning, Low-Rank, ICML
11footnotetext: University of Toronto and 2Vector Institute. Equal contribution. Correspondence to: Stephen Zhang $<$stephenn.zhang@mail.utoronto.ca$>$ and Mustafa Khan $<$mr.khan@mail.utoronto.ca$>$.
@@ 1 Introduction
1.1 ‘Catch, Tag, and Release’ in Aquatic Conservation
In marine biology, the ‘catch, tag, and release’ mechanism is a vital tool for tracking fish populations. A fish is caught, fitted with a tracking tag encoding critical information, and then released back into a water stream, where it is monitored over time. This enables researchers to understand migration patterns, survival rates, and ecosystem interactions.
1.2 Pervasive Phenomena in LLMs
Remarkably, modern large language models (LLMs) exhibit a strikingly similar mechanism when processing tokens. To understand why, we first examine two recent discoveries in LLM research:{mdframed}[shadow=true, shadowsize=1pt] Attention Sinks: Coined by Xiao et al. (2024), the authors found that all tokens attend strongly to the first token. Follow-up works demonstrated that sinks may also appear in later tokens (Yu et al., 2024; Sun et al., 2024a; Cancedda, 2024).
{mdframed}
[shadow=true, shadowsize=1pt] Outlier Features: Feature dimensions in the activations that are significantly larger in magnitude (Kovaleva et al., 2021; Dettmers et al., 2022). Unlike attention sinks, outlier feature dimensions are consistent across tokens.
(a) Catch: Attention sinks catch the attention of a subset of tokens, through outlier features, as shown in Figures 2(a) and 2(b).
(b) Tag: The attention layer then tags the captured tokens by introducing an outlier feature dimension across all tokens. This shared tag causes tokens to cluster based on the attended sink, e.g. in PCA plots, as shown in Figure 2(c).
(c) Release: The tagged tokens are then released into the residual stream, where deeper layers can exploit the tag for further processing. This common tag induces token clustering even in deeper layers, e.g. in PCA plots, as shown in Figure 2(d).
@@ 1.2 Pervasive Phenomena in LLMs
Both phenomena have been linked to model performance, particularly in streaming (Xiao et al., 2024; Guo et al., 2024c), quantization techniques (Dettmers et al., 2022; Son et al., 2024; Ashkboos et al., 2024), and pruning techniques (Sun et al., 2024b; Zhang and Papyan, 2024), as behaviors that need to be preserved in the model’s attention weights and features. This, in turn, has driven prior research to investigate the following question.
1.3 Why Do Sinks and Outliers Emerge?
Existing answers generally fall into two main categories:
To Create Implicit Attention Biases
Models lack explicit bias parameters in the attention layers, leading attention sinks and large outliers to emerge as compensatory mechanisms that artificially introduce such biases. As a result, these effects can be mitigated by incorporating explicit key or value bias parameters (Sun et al., 2024a; Gu et al., 2024).
To Turn Off Attention Heads
Attention sinks emerge as a learned, data-dependent mechanism that disables an attention head when it is not needed for the current sequence (Bondarenko et al., 2023; Guo et al., 2024b).
@@ 1.4 Open Questions
Both perspectives still leave the following unanswered questions:
@@
    How do attention sinks and outlier features manifest in model parameters?
While sinks and outlier features play a crucial role in model compression, prior research has primarily focused on their presence in model embeddings. Their manifestation within model parameters remains largely unexplored, raising questions about the best strategies for preserving these critical properties during compression.
@@
    Why do attention sinks emerge in later tokens?
Neither perspective fully explains why attention sinks appear in later tokens alongside the first token sink (Yu et al., 2024; Cancedda, 2024). These sinks are sequence-dependent, contradicting the bias perspective, and the presence of multiple sinks seems unnecessary from the active-dormant perspective.
@@
    Are attention sinks beneficial or better dispersed?
It remains unclear whether attention sinks are critical for model performance and need to be preserved (Xiao et al., 2024; Son et al., 2024), or if they should be dispersed or removed altogether (Sun et al., 2024a; Yu et al., 2024).
(a) Attention Weights. Two distinct attention sinks catch the attention of tokens appearing later in the sequence.
(b) Output of Attention Layer. The sink token exhibits an outlier feature that is then copied by the attention weights to the caught tokens, creating an outlier feature dimension.
@@
(d) PCA on the residual stream in a deep layer. Tagged tokens propagate through the residual stream, clustering based on their previously attended sink. Tokens that attended to the first sink are green, while those attending to the second sink are brown and yellow.
Figure 2: Visualization of the ‘catch, tag, and release’ mechanism. The second and third subplots use PCA-based coloring of embeddings.
@@ 1.5 Contributions
Our first contribution is a theoretical study of a two-layer attention model trained to average a sequence of numbers. This analysis answers the questions posed and establishes the following claims, summarized in the schematic below:
Low-Rank Structure in ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V}$ Attention Sinks and Outlier Features ‘Catch, Tag, and Release’ Mechanism Sequence Averaging, Few-Shot Learning, …
@@
    A low-rank structure in the attention weight matrices, ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V}$, is responsible for attention sinks and outlier features.
@@
    Attention sinks and outlier features are leveraged by the model to simulate a ‘catch, tag, and release’ mechanism, depicted and explained in Figure 1.
@@
    The mechanism is useful for averaging a sequence, few-shot learning, and many other tasks.
Our second contribution involves showing that pruning algorithms that do not preserve low-rank structure in weight matrices, destroy attention sinks and feature outliers, harm the ‘catch, tag, release’ mechanism, and perform worse on few-shot learning.
1.6 Relationship with Prior Perspectives
The ‘catch, tag, and release’ mechanism does not contradict the perspectives on attention sinks presented in prior works by Sun et al. (2024a); Gu et al. (2024); Guo et al. (2024b). We hypothesize that the prior perspectives are accurately describing the behavior of the attention sink associated specifically with the first token in the sequence. The ‘catch, tag, and release’ mechanism expands on this by describing the behavior associated with the data-dependent sinks that emerge in later tokens (Yu et al., 2024).
2 Evidence for ‘Catch, Tag, Release’
This subsection provides empirical evidence for the existence of the ‘catch, tag, release’ mechanism. To this end, we input a prompt into the Phi-3 Medium model (Abdin et al., 2024), and compute various statistics, each offering evidence for the existence of a specific component of the mechanism.
@@ 2 Evidence for ‘Catch, Tag, Release’
As described in Figure 1(a), the evidence of the catch mechanism amounts to showing the existence of an attention sink. We therefore save the attention weights of layer $3$, head $6$ and plot them in Figure 2(a). The visualization shows that there are indeed two sinks that are catching the attention of subsequent tokens.
@@
As described in Figure 1(b), demonstrating the tagging mechanism requires verifying token clustering based on the attended sink. Following Oquab et al. (2024), we compute the top two or three principal components of the $d\times d$ covariance matrix of the attention head outputs, where $d$ is the embedding dimension, and project the embeddings onto this basis. The projected values are then normalized to $[0,255]$ and mapped to RGB channels. Figure 2(c) illustrates the results, showing a clear grouping of tokens by their attention sinks.
@@
As described in Figure 1(c), evidence for release amounts to showing that tokens in the residual stream in deeper layers have the same clustering behavior as exhibited in the attention head output. We therefore hook the inputs into the feedforward network in layer $17$, and apply the same PCA projection and normalization step as described earlier. The results, presented in Figure 2(d), reveal a similar grouping, providing evidence that the model has utilized the tags generated in the earlier layer.
Appendix B presents additional measurements using various prompts including chain-of-thought (Wei et al., 2022), zero-shot chain-of-thought (Brown et al., 2020), among others.
3 Intervention Through Model Pruning
This subsection aims to intervene at the root of the logical sequence, displayed in Section 1.5, namely the low-rank structure of the attention weight matrices and to test its effect on the attention sinks and outlier features, the ‘catch, tag, release’ mechanism, and ultimately performance on a down-stream task which relies on it: few-shot learning. The intervention is done through the use of pruning algorithms.
3.1 How is Pruning Related?
Pruning algorithms compress models by retaining the fewest possible nonzero parameters while preserving performance (Mozer and Smolensky, 1988; LeCun et al., 1989; Hassibi and Stork, 1992). However, they typically do not prioritize accurately capturing low-rank structures in the weights.
An exception to this is a recently proposed compression algorithm for LLMs, called OATS (Zhang and Papyan, 2024), which approximates the model’s weight matrices through a sparse-plus-low-rank decomposition:
${\bm{W}}\approx\underbrace{{\bm{S}}}_{\text{Sparse}}+\underbrace{{\bm{U}}{\bm{V}}}_{\text{Low-Rank}}\!\!\!\!\mathrlap{{}^{\top}}\ \ .$
This method, by design, explicitly preserves low-rank structures through the low-rank term, a key feature absent from traditional pruning algorithms.
Comparing pruning algorithms with OATS is therefore a natural intervention since it directly controls for the low-rank structure. Having established the intervention, we now test its effect on the attention sinks and outlier features.
3.2 Low-Rank Terms $\implies$ Sinks and Outliers
Figure 4 presents the attention weights for layer $2$, attention head $6$, of a Phi-3 Medium model in three configurations: a dense model, a model compressed by $50$% using OATS, and a model compressed by $50$% using OATS where the low-rank matrices for ${\bm{W}}_{Q},{\bm{W}}_{K},{\bm{W}}_{V}$ for that layer are set to $\bm{0}$. The plots reveal clear sinks and outliers, except in the model lacking low-rank terms, suggesting that the low-rank term in OATS is indeed responsible for both phenomena.
3.3 No Low-Rank Terms $\implies$ No Sinks and Outliers
Figure 5 compares the attention weights in layer $36$, head $16$, of a dense Phi-3 Medium model with models that are pruned by $50$% using OATS, and pruned by $50$% using Wanda (Sun et al., 2024b), and SparseGPT (Frantar and Alistarh, 2023) – two methods that do not prioritize a low-rank structure in their formulation. The visualizations reveal that pruning methods like Wanda and SparseGPT lack certain attention sinks found in both the dense and OATS models.
3.4 Sinks and Outliers $\implies$ ‘Catch, Tag, and Release’
The previous two subsections show that classical pruning algorithms fail to retain the low-rank components of attention weights, which results in the loss of attention sinks and outlier features. In contrast, OATS preserves them. Building on Section 2, which establishes the necessity of sinks and outliers for the ‘catch, tag, release’ mechanism, we conclude that pruning algorithms like Wanda and SparseGPT break the ‘catch, tag, release’ mechanism, whereas OATS maintains it.
3.5 ‘Catch, Tag, and Release’ $\implies$ Few-Shot Learning
Since the ‘catch, tag, release’ mechanism segments sequences, it is likely crucial for distinguishing between examples in few-shot learning. To test this, we measure the gap in $k$-shot performance on the MMLU dataset (Hendrycks et al., 2021) using LM Harness (Gao et al., 2024), where $k$ ranges from 00 to $5$. We compare the performance of OATS with Wanda and SparseGPT in Figure 3.

### v2

# Attention Sinks: A ‘Catch, Tag, Release’ Mechanism for Embeddings

@@ Abstract
Large language models (LLMs) often concentrate their attention on a few specific tokens referred to as attention sinks. Common examples include the first token, a prompt-independent sink, and punctuation tokens, which are prompt-dependent. While the tokens causing the sinks often lack direct semantic meaning, the presence of the sinks is critical for model performance, particularly under model compression and KV-caching. Despite their ubiquity, the function, semantic role, and origin of attention sinks—especially those beyond the first token—remain poorly understood. In this work, we conduct a comprehensive investigation demonstrating that attention sinks: catch a sequence of tokens, tag them using a common direction in embedding space, and release them back into the residual stream, where tokens are later retrieved based on the tags they have acquired. Probing experiments reveal these tags carry semantically meaningful information, such as the truth of a statement. These findings extend to reasoning models, where the mechanism spans more heads and explains greater variance in embeddings, or recent models with query-key normalization, where sinks remain just as prevalent. To encourage future theoretical analysis, we introduce a minimal problem which can be solved through the ‘catch, tag, release’ mechanism, and where it emerges through training.
@@ 1 Introduction
1.1 ‘Catch, Tag, Release’ in Aquatic Conservation
In marine biology, the ‘catch, tag, release’ mechanism is a vital tool for tracking fish populations. A fish is caught, fitted with a tracking tag encoding critical information, and then released back into the water stream, where it can be monitored over time. This process enables researchers to understand migration patterns and ecosystem interactions. Remarkably, LLMs exhibit a strikingly similar mechanism when processing tokens. To understand why, we first examine a recent discovery in LLM research.
1.2 The Prevalence of Attention Sinks
In many models, tokens often focus disproportionately on a select few positions in the sequence. Xiao et al. (2024) identified the first token as a common focal point, referring to it as an attention sink. Follow-up studies have shown that, beyond this fixed sink, additional ones can emerge based on the input – often appearing on punctuation (Yu et al., 2024; Sun et al., 2024; Cancedda, 2024).
Preserving these sinks has proven critical for retaining model performance in several key areas, including KV-caching (Xiao et al., 2024; Liu et al., 2024; Guo et al., 2024b; Willette et al., 2024), quantization (Lin et al., 2024; Son et al., 2024), and pruning (Zhang and Papyan, 2025a). Motivated by these findings, researchers have begun to investigate the following question.
(a) Catch: Each box corresponds to a different token at the input of the attention layer, whose activation is denoted by ${\bm{e}}_{i}$, and the arrows represent attention interactions. The attention sinks ${\bm{e}}_{\text{sink}_{1}}$ and ${\bm{e}}_{\text{sink}_{2}}$ catch the attention of tokens ${\bm{e}}_{2},{\bm{e}}_{3},\ldots,{\bm{e}}_{t-1}$ and ${\bm{e}}_{t+1},{\bm{e}}_{t+2},\ldots,{\bm{e}}_{T}$, respectively. This causes vertical bands to emerge in the attention weights $\bm{A}$}, as shown in Figure 2(a).
(b) Tag: The left grid shows the attention value matrix ${\bm{V}}=[{\bm{e}}_{1};\ldots;{\bm{e}}_{T}]$, where activation vectors ${\bm{e}}_{i}$ are stacked vertically. The right grid shows the output of the attention layer ${\bm{Z}}=[{\bm{z}}_{\text{sink}_{1}};{\bm{z}}_{2};\ldots;{\bm{z}}_{T}]={\bm{A}}{\bm{V}}$, with output vectors ${\bm{z}}_{i}$ also stacked vertically. The value vectors of the sinks, ${\bm{e}}_{\text{sink}1}$ and ${\bm{e}}_{\text{sink}_{2}}$, are copied to all tokens that attend to them, thereby tagging them. These tags cause the token representation to cluster based on the sink they attended to, as revealed in the PCA plot in Figure 2(c). The inputs to the attention layer, prior to the tagging, show no such clustering, as shown in Figure 2(b).
(c) Release: Each box corresponds to a different token at the output of the attention layer. The attention outputs are added to the residual stream as ${\bm{e}}_{i}+{\bm{z}}_{\text{sink}}$, creating common directions in representation space, in the form of the tags ${\bm{z}}_{\text{sink}_{1}}$ and ${\bm{z}}_{\text{sink}_{2}}$ shared across multiple tokens. These tags cause the token representations to cluster in deeper layers, as revealed in the PCA plot in Figure 2(d).
@@ 1.2 The Prevalence of Attention Sinks
1.3 Why Do Attention Sinks Emerge?
Existing answers generally fall into three main categories:
To Create Implicit Attention Biases:
Attention layers lack explicit bias parameters. Attention sinks emerge as compensatory mechanisms that artificially introduce such biases, and they can be mitigated by incorporating explicit key or value bias parameters (Sun et al., 2024; Darcet et al., 2024; Gu et al., 2024).
To Turn Off Attention Heads:
Attention heads are not needed for certain sequences. Attention sinks emerge as a learned, data-dependent mechanism that effectively disables them by capturing nearly all the attention (Bondarenko et al., 2023; Guo et al., 2024a).
To Prevent Over-Mixing of Tokens:
Transformers are prone to excessive token mixing and rank collapse (Geshkovski et al., 2024; Barbero et al., 2024, 2025b). First-token attention sinks help mitigate these issues by anchoring the sequence and limiting uncontrolled interaction across positions.
@@ 1.4 Open Questions
While each of the three perspectives shine a light on the role of first-token attention sink, they still leave the following questions unanswered:
@@
    Why do attention sinks emerge in later tokens?
None of the perspectives addresses this question. These sinks are sequence-dependent (Yu et al., 2024; Cancedda, 2024), contradicting the bias perspective, and the presence of multiple sinks seems unnecessary from both the active-dormant and over-mixing perspective.
@@
    Do the representations of attention sinks—despite corresponding to semantically meaningless tokens like punctuation—nonetheless encode meaningful information?
This question remains largely unexplored (Yu et al., 2024), leading to conflicting lines of research, with some studies exploring how sinks can be preserved (Xiao et al., 2024; Son et al., 2024), while others investigating how they can be dispersed or removed (Sun et al., 2024; Yu et al., 2024; Zuhri et al., 2025; Fu et al., 2025; Kang et al., 2025).
@@
    Are there differences in the number of attention sinks between pretrained LLMs and those fine-tuned for reasoning tasks (DeepSeek-AI, 2025)?
If attention sinks play a functional role in reasoning, one would expect them to be more prevalent in models optimized for such purposes. Furthermore, the locations of these sinks may align with semantically meaningful boundaries, segmenting content in ways that support structured reasoning.
  • Q4:
    How does query-key normalization affect attention sinks?
It is conceivable that query-key (QK) normalization (Henry et al., 2020) alters the effective temperature of the softmax distribution in attention layers. This, in turn, may smoothen or sharpen attention weights, potentially affecting the formation of attention sinks.
(a) Attention Weights. Two attention sinks catch the attention of subsequent tokens in the sequence.
(b) PCA on input to the attention layer. Tokens exhibit no clustering.
@@
(d) PCA on the residual stream in a deeper layer. Tagged tokens propagate through the residual stream, clustering in a deeper layer based on their previously attended sink. Tokens that attended to the first sink are green, while those attending to the second sink are brown and yellow.
Figure 2: Qualitative analysis of the ‘catch, tag, release’ mechanism. The second and third subplots use PCA-based coloring of embeddings, described in Section 2. Appendix A presents additional measurements across a wide range of models, layers, attention heads, and prompts, including chain-of-thought (Wei et al., 2022), and zero-shot chain-of-thought (Brown et al., 2020).
@@ 1.5 Contributions
Through an empirical study, we answer the questions posed above and establish the following claims:
@@
    Attention sinks implement a ‘catch, tag, release’ mechanism, the steps of which are detailed in Figure 1 (Sections 2, 3).
@@
    Probing experiments reveal that the resulting tags are not arbitrary – they encode semantically meaningful information, such as the truth value of a statement (Section 4).
@@
    Compared to pretrained models, DeepSeek-distilled models exhibit the mechanism across more attention heads and explain a greater proportion of variance in the embeddings, indicating a stronger and more pronounced instantiation of the mechanism (Section 5).
  • A4:
    Attention sinks remain prevalent in models with QK normalization, despite the normalization explicitly imposed on tokens prior to computing attention scores. (Section 6).
To support future theoretical analysis, we introduce a minimal problem that is solvable via the explicit use of the ‘catch, tag, release’ mechanism and demonstrate empirically that the mechanism naturally emerges through standard training.
2 Visualizing the ‘Catch, Tag, Release’ Mechanism
This section presents a visual exploration of the ‘catch, tag, release’ mechanism. Figure 2 illustrates a representative example selected for clarity, but similar behavior consistently emerges across a wide range of models, prompts, layers, and attention heads. To support this generality, we include additional visualizations in Appendix A.
@@ 2 Visualizing the ‘Catch, Tag, Release’ Mechanism
As illustrated in Figure 1(a), the evidence of the catch mechanism amounts to showing the existence of an attention sink. We therefore feed a prompt to the Phi-3 Medium model (Abdin et al., 2024a), and save the attention weights of layer $3$, head $6$ and plot them in Figure 2(a). The visualization shows that there are two sinks that are catching the attention of subsequent tokens.
@@
As shown in Figure 1(b), demonstrating the tagging mechanism requires verifying token clustering based on the attended sink. Following Oquab et al. (2024), we compute the top two or three (depending on the setting) principal components of the $d\times d$ covariance matrix of the attention head outputs, where $d$ is the embedding dimension, and project the activations onto this basis. The projected values are then normalized to $[0,255]$ and mapped to RG or RGB channels. Figure 2(c) illustrates the results, showing a clear grouping of tokens by their attention sinks at the output of the attention head. In contrast, Figure 2(b) depicts the absence of such grouping at the input to the attention head.
@@
As depicted in Figure 1(c), evidence for release amounts to showing that tokens in the residual stream in deeper layers have the same clustering behavior as exhibited in an earlier attention head output. We therefore hook the inputs into the feedforward network in layer $17$, and apply the same PCA projection and normalization step as described earlier. The results, presented in Figure 2(d), reveal a similar grouping, providing evidence that the model has utilized the tags generated in the earlier layers to cluster the embeddings.
3 Measuring the ‘Catch, Tag, Release’ Mechanism
We provide a quantitative analysis to substantiate the presence of the ‘catch, tag, release’ mechanism across a wide range of model families, including Qwen 2.5 (Yang et al., 2024), Phi-3 (Abdin et al., 2024a), LLaMA-3 (Grattafiori et al., 2024), and Mistral 7B (Jiang et al., 2023). The analysis is aggregated over 170 prompts collected by Gu et al. (2024), truncated to 150 tokens.
3.1 Identifying Attention Sinks and Their Associated Tags
We leverage the metric proposed by Gu et al. (2024) to identify which tokens are attention sinks. Letting $\bm{A}$} denote the attention weights (i.e. softmax probabilities) of an attention head, token $t$ is identified as an attention sink for that head if and only if:
$\alpha_{t}:=\frac{1}{T-t+1}\sum_{k=1}^{T}{\bm{A}}_{k,t}>\epsilon,$ (1)
where $T$ is the length of the prompt and $\epsilon$ is a predetermined threshold that is set to $\epsilon=0.2$ for all the following experiments [^1] A sensitivity analysis of the threshold choice is provided in Appendix F..
For any token $t$ designated as an attention sink, its tag is defined to be its value vector extracted from the attention head’s value matrix: ${\bm{v}}_{t}={\bm{V}}_{t,\;:}\in\mathbb{R}^{d_{\text{head}}}$, where $d_{\text{head}}$ is the head dimension.
3.2 Quantifying Variance Explained by Tags
Suppose $n$ tokens are identified as attention sinks. We concatenate their tags into a matrix $\mathbf{V}_{\text{tag}}\in\mathbb{R}^{n\times d_{\text{head}}}$ and apply Principal Component Analysis (PCA):
$\mathbf{V}_{\text{tag}}^{\top}\mathbf{V}_{\text{tag}}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top},$
where $\bm{U}$} is the matrix of eigenvectors and $\bm{D}$} is the diagonal matrix of eigenvalues. The matrix $\mathbf{U}\mathbf{U}^{\top}$, which has rank $n$, defines a projection onto the subspace spanned by the tags. We take the output of the attention head, given by ${\bm{A}}{\bm{V}}$ (where $\bm{A}$} and $\bm{V}$} denote the attention weights and values, respectively), and project it onto this subspace, yielding the:
$\text{Variance Explained}=\frac{\|\mathbf{A}\mathbf{V}\mathbf{U}\mathbf{U}^{\top}\|_{F}}{\|\mathbf{A}\mathbf{V}\|_{F}}.$ (2)
This quantity is upper bounded by one with equality only if the attention head’s output lies entirely within the tag subspace, providing a soft measure of how much of the output is explained by the tags.
Figure 4 below shows that 1-2 tags typically explain $20\%-40\%$ of the variance in the outputs of the attention heads, but in many cases up to $70\%$. This provides evidence that the tokens are being tagged and that these tags contribute significantly to the tokens’ representations.
     Qwen Family
      Phi Family
    LLaMA Family
Mistral
(a) Attention Sink Counts. Counts are computed for each head using Equation 1, and their distribution is visualized for each model using violin plots.
     Qwen Family
      Phi Family
    LLaMA Family
Mistral
(b) Variance Explained by Tags. The metric is computed for each head using Equation 2, and summary statistics are shown for each model using box plots. Boxes indicate the $25\text{th}$, $50\text{th}$, and $75\text{th}$ percentiles, while whiskers represent the $5\text{th}$ and $95\text{th}$ percentiles.
Figure 4: Quantitative analysis of the ‘Catch, Tag, Release’ mechanism.
4 Investigating the Semantic Role of Tags
The previous section demonstrated that attention sinks tag tokens. This observation prompts a deeper question: do the distributed tags carry semantically meaningful information that is being released back into the residual stream as part of the release mechanism? To answer this question, we turn to a dataset designed to test semantic encoding.
4.1 Cities Dataset (Marks and Tegmark, 2024)
The city of Tokyo is in Japan. This statement is: TRUE. The city of Hanoi is in Poland. This statement is: FALSE. The city of [CITY] is in [COUNTRY]. This statement is:
Figure 5: cities prompt example.
The dataset consists of prompts involving ([CITY], [COUNTRY]) pairs that render a statement either true or false (see Figure 5). The goal is to assess whether the activations of individual tokens can be used to classify the True/False label of the statement. We focus specifically on the final period token – located at position $t$, indicated by the arrow in Figure 5 – as punctuation tokens frequently cause attention sinks.
4.2 Decomposing Activations into Tag and Non-Tag Components
The decomposition of the activation is given by:
${\bm{z}}=\sum_{k=1}^{T}{\bm{A}}_{t,k}{\bm{V}}_{k,\;:\;}={\bm{z}}_{\text{tag}}+{\bm{z}}_{\text{no tag}}$
where: